OpenStudy (anonymous):

2^(-6x)=15^(x-3) Solve for x . Write the exact answer using base-10 logarithms.

OpenStudy (anonymous):

ok actually it is \[-6x\log(2)=(x-3)\log(5)\]

OpenStudy (anonymous):

multiply out on the left get \[-6x\log(2)=x\log(5)-3\log(5)\] then \[x\log(5)+6x\log(2)=3\log(5)\] \[x\left(\log(5)+6\log(2)\right)=3\log(5)\] \[x=\frac{3\log(5)}{\log(5)+6\log(2)}\]

OpenStudy (anonymous):

thats the right answer

OpenStudy (anonymous):

except everywhere i wrote a 5, you should write a 15

OpenStudy (anonymous):

satellite can you come help me please!!!!!!!!!! after this